# Regimes of Steady-State Turbulence in a Quantum Fluid

### v1

@@ I Introduction
A cascade of energy between scales can emerge when energy is injected into a fluid at a particular length scale and dissipated at another scale. This results in the formation of ever smaller structures down to the dissipation scale (direct cascade) [4], or energy accumulating into coherent structures at the system scale (inverse cascade) [17]. If the interaction is local in $k$-space, power laws can appear in the energy distribution associated with the energy cascade, providing clear signatures of particular dynamics, which may otherwise prove opaque. The most well-known example of emergent power-law behavior in fluid turbulence is the Kolmogorov-Obhukov $-5/3$ law [4]. This states that at high Reynolds number in classical incompressible fluids, the distribution of kinetic energy in Fourier-space is given by $E(k)=C\varepsilon^{2/3}k^{-5/3}$, where $C$ is a universal dimensionless constant and $\varepsilon$ is the rate of energy flux per unit mass. Recent experiments [5, 9, 15] and numerical investigations [18, 19] of turbulent BECs have reported power-laws related to predictions from weak-wave turbulence (WWT) theory [20, 21], observed in the wave-action spectrum accessible in experiments.
@@
For weak forcing, we observe $k^{-3.5}$ scaling as reported in experiments and other numerical studies. Further decomposition shows that the compressible and quantum pressure components are in lock step throughout the power-law range of $k$, consistent with a Bogoliubov-phonon direct cascade. For strong forcing, the system contains many vortices, and approaches a $k^{-7/3}$ power law for the compressible kinetic energy only, consistent with the WWT prediction of inverse particle cascade scaling in the non-condensed regime [19]. The phase coherence length of the GPE wave function drops to the healing length in this regime associated with dense vortex tangles.
@@ II Theoretical Model
Under conditions of weak forcing, the GPE provides a quantitative treatment of the system, directly comparable with experiments. Stronger forcing involves large energy injection and requires careful consideration. Due to the numerical scale of the system and the challenges of running large-scale simulations for very long physical times (of order 10s), in this work we take the approach of constructing a minimal GPE model, similar to other recent numerical treatments [5, 23]. Our primary differences are using higher working precision which allows consideration of stronger forcing, and larger spatial sizes in units of the healing length. Ideally, we aim to observe a power-law over a decade of wavenumbers in order to rule out the possibility of falsely attributing power-law behavior to a crossover [34].
@@
where $g=4\pi\hbar^{2}a/m$, $\hbar$ is the reduced Planck constant, $a$ is the $S$-wave scattering length, and $m$ is the atomic mass. For a homogeneous ground state with chemical potential $\mu$, we define the healing length $\xi$ via
@@
to coordinates $x-L_{x}$, etc, where $H(x)$ is the Heaviside step. Over a region of width $\ell_{b}=6\xi$ centered on $x=3\xi$ the potential $V_{\textrm{wall}}(x)$ ramps from zero up to $V_{b}=30\mu$, providing sufficiently steep walls for box-confined dynamics while avoiding overly sharp changes in the potential; the shape of the potential causes the flat density to drop away steeply at $x=0$, so that after the shift, the homogeneous region is constrained to ±$\pm L_{x}$. Numerically finding the ground state $\psi_{0}(\mathbf{r})$ of $V_{\textrm{static}}(\mathbf{r})$ using imaginary time evolution for chemical potential $\mu=k_{B}\times 1$nK, results in a total atom number $N\simeq 7.2\times 10^{5}$. The loss of initial anisotropy of the ground state under time evolution provides provide a useful indicator of the development of steady isotropic turbulence.
@@
The momentum required for a particle to escape is a function of the height of the trap walls [15], and for our trapping potential loss occurs at dimensionless dissipation wavenumber $k_{D}\xi=\sqrt{2V_{b}/\mu}=\sqrt{60}\simeq 7.74$  [^1] This choice also gives sufficient margin between the dissipation $k$-value and the Nyquist $k$-values for a grid of $256^{3}$ points (Appendix A). High-energy particles are thus dissipated well before their spatial grid representation becomes inaccurate. . This energy and spatially selective loss is the only dissipation used in our simulations, and provides the energy sink required for a steady state to form under conditions of steady forcing.
To create turbulence in the forced and damped GPE, we apply $V_{F}$ to the ground state $\psi_{0}$ and evolve until $t=t_{f}\equiv 10$ seconds. This time interval ensures a steady state is reached for all $U_{F}$ values. We use a grid of $256^{3}$ points and time-evolve using a pseudo-spectral Runge-Kutta method, referred to as the RK4IP [37]. A fixed time step $\Delta t=10^{-3}$s is chosen to ensure numerical convergence all simulations. The code was written in Julia and optimized to achieve a minimal memory footprint and run on GPU hardware using the CUDA.jl library [38, 39]. We achieved a runtime of 38 minutes per second of dynamics on NVIDIA A100 GPU using Float64 representation. For details of the algorithm and error analysis see Appendix A.
Figure 1: Time series of density isosurfaces of the cuboid box trap shaken along $z$ with forcing amplitude $U_{F}=\mu$: (a) $t=0.0$s shows the initial ground state of the box trap; (b) $t=0.156$s weak density excitations along $z$; (c) $t=0.391$s nucleation of vortex rings; (d) $t=10.0$s steady state turbulence. Vortex lines can be seen in red. In the steady state there are many small vortex rings, but no long vortex lines.
@@ III.1 Vortices and density waves
To outline the general phenomenology, we first present the dynamics that occur at a forcing amplitude of $U_{F}=\mu$. The initial ground state, shown in Fig. 1(a), responds to forcing by weakly oscillating between the sides of the trap, creating weak density waves, as seen in Fig. 1(b). After sufficient energy is injected, the superfluid develops many small vortex rings parallel to the $x-y$ plane [Fig. 1(c)]. After approximately two seconds, the system transitions to an isotropic density and vortex distribution involving many vortices and significant energy, density, and phase fluctuations [Fig. 1(d)]. For this particular forcing, there are no extended vortex lines in the steady state, but many small vortex rings are present.
@@
Fig. 3(a) and (b) show the normalization and energy per particle of the system during excitation. We observe a brief delay between when forcing begins and when particles have sufficient energy to escape the trap, shown in Fig. 3(a). Consequently, the energy per particle of the system rapidly increases [Fig. 3(b)]. Once dissipation takes effect, the energy per particle approaches an approximate steady-state, oscillating about a mean value that increases with $U_{F}$; the system thus approaches an energetic steady state while continuing to lose particles.
@@ III.2 Energy decomposition
$E_{\text{kin}}$ $=\frac{m}{2}\int d^{3}\mathbf{r}|\mathbf{u}^{\alpha}(\mathbf{r})|^{2}.$ (13)
Figure 3: Development of turbulence (a) Atom number over time. The rate of particles escaping the trap increases with $U_{F}$. (b) Total energy per particle over time. (c), (d) Total incompressible and compressible kinetic energy during initial period of excitation. (e) Kinetic energy fractions in the steady state ($t=10.0$s) as a function of $U_{F}$. Dashed lines show the four special cases of Fig. 2.
Figures 3(c) and (d) show $E_{\textrm{kin}}^{i}$ and $E_{\textrm{kin}}^{c}$ during the initial period of excitation. Both components oscillate in response to the forcing, acquiring a mean value that increases with $U_{F}$. The kinetic energy components of the steady-state display a nonlinear dependence on $U_{F}$, indicating different regimes of turbulence, shown in Fig. 3(e). The relative importance of compressible and incompressible energy indicates a change from wave to vortex turbulence. Significant kinetic energy is stored in density gradients even in the absence of vortices, evident in the quantum pressure.
Figure 3(e) is consistent with the behavior shown in Fig. 2: from $U_{F}=0.2\mu$, $E^{c}_{\textrm{kin}}$ sharply increases as density fluctuations grow [Fig. 3(e)]; for the same $U_{F}$ values, $E^{i}_{\textrm{kin}}$ increases gradually as vortices are present only at the condensate surface. Recent reports of wave-turbulence in BECs generally correspond to these forcing amplitudes ($U_{F}<\mu$) [5, 15, 19, 9], which is consistent with the dominance of compressible kinetic energy in this region.
At $U_{F}\simeq 1.2\mu$, a distinct change is evident in Fig. 3(e). $E^{i}_{\textrm{kin}}$ increases rapidly, associated with the formation of bulk vortices, and $E^{c}_{\textrm{kin}}$ has a much weaker response to increased forcing. Finally, another change in response occurs for $U_{F}>2\mu$, where incompressible kinetic energy and quantum pressure have a weaker response to increased forcing. In Fig. 2 this corresponds to the steady state vortex distribution becoming homogeneous. There may be a limit to the number of vortices that can stably exist within a BEC of a certain size and density, bounding growth of incompressible kinetic energy with forcing.
This analysis suggests distinct turbulent regimes exist for different forcing strengths. In the following sections, we apply tools of spectral analysis to relate these regimes to previous observations and theoretical predictions, aiming to identify power law behavior associated with particular regimes in Fig. 3(e).
@@ IV Steady State turbulence
Figure 5: (a)-(c) Decomposed momentum distributions for representative forcing amplitudes of Fig. 2, averaged over one forcing period. (a) As forcing increases a $k^{-3.5}$ scaling regime is evident in both compressible and quantum pressure components. (b) Compressible and quantum pressure components decouple, associated with bulk vortices. (c) For very strong forcing, the compressible wave-action spectrum approaches a $k^{-7/3}$ power-law form. (d) Ratio of total incompressible to compressible kinetic energy, with identified power law regimes. $U_{F}$ values from Fig. 2(b) are shown as dashed lines.
@@
We do not observe any obvious power-law behaviour for $U_{F}\ll\mu$ (not shown). Within a narrow region of forcing, $0.15\mu\lesssim U_{F}\lesssim 0.3\mu$, the compressible density approximates $k^{-3}$. However, the approximate scaling only extends over a factor of 2 in wave number, insufficient to identify power-law behaviour [34]. Starting at $U_{F}=0.4\mu$, we observe a clear $k^{-3.5}$ scaling over nearly a decade of $k\xi$, in both compressible and quantum pressure momentum distributions, shown in Fig. 5(a). This power-law remains intact for a wide interval of forcing energies, $0.4\mu\leq U_{F}\leq 1.2\mu$, and as incompressible energy remains small, the power-law is also evident in the total spectrum $n(k)$. This is consistent with the observed vortex configurations [Fig. 2(a)], where extended bulk vortices remain rare. In this regime it is notable that $n^{c}(k)\simeq n^{q}(k)$ in the power law region of wavenumbers. We return to this point in Sec. V.1.
@@
For $U_{F}\gtrsim 1.2\mu$, extended vortices develop in the bulk superfluid in the steady-state, and the system enters a transitional regime without power-law behavior in any component of kinetic energy [Fig. 5(b)]. The total wave-action spectrum gradually steepens in the ultraviolet due to an increase in the incompressible energy-density, while the compressible and quantum pressure components uncouple, and the slope of $n^{c}(k)$ decreases. The appearance of bulk vortices is associated with the breakdown of $k^{-3.5}$ scaling. Eventually, we observe a wide $n^{c}(k)\propto k^{-7/3}$ power-law for strong forcing, $U_{F}\gtrsim 2\mu$, in agreement with the predicted inverse particle cascade for weak forcing in the absence of a condensate [Fig. 5(c)]  [19].
@@
To further understand the different regimes of steady turbulence, in Fig. 5(d) we plot the ratio of incompressible to compressible kinetic energy, with the identified power-law regimes. There is a clear change in the energy dependence on $U_{F}$ at the boundary of each power-law regime. Initially, the ratio is rapidly decreasing, due to compressible excitations paired with a lack of vortex nucleation. With additional energy injected, this is succeeded by a sharp increase in the kinetic energy ratio (as surface vortices begin to proliferate), and the appearance of a $k^{-3.5}$ power-law. The breakdown of $k^{-3.5}$ involves extended vortices entering the bulk in the steady-state, and a steep rise in incompressible energy with $U_{F}$. The system enters a regime of vortex turbulence, but shows no scaling behavior. Finally, for $U_{f}\gtrsim 2\mu$, a $k^{-7/3}$ scaling emerges where vortices are approximately homogeneous and dense, and the incompressible energy responds only weakly to increased forcing. In this regime $E^{i}_{\text{kin}}/E^{c}_{\text{kin}}>1$, a regime we can regard as strong quantum vortex turbulence, in contrast to weak wave turublence where $E^{i}_{\text{kin}}/E^{c}_{\text{kin}}\ll 1$.
While Fourier-space reveals power-law behavior, we can also examine the corresponding correlation functions in position-space. When power-law behavior weakens, the appearance of characteristic length scales can be more clearly understood in position space. Fourier-transformation of the momentum distribution gives the system-averaged spatial two-point correlation of the Gross-Piteavskii wave function, as discussed in Sec. III.3. In Fig. 6(a), we plot the system averaged two-point correlation function $g_{1}(r)\equiv G_{\psi\psi}(r)/G_{\psi\psi}(0)$, calculated by transforming $\langle\psi\|\psi\rangle(k)=4\pi k^{2}n(k)$ to position space via (20). The two-point correlator of the ground state ($U_{F}=0$) decays over a scale comparable to the shortest cuboid length $L_{z}$. As $U_{F}$ increases, the length scale is reduced. Eventually, for strong forcing the correlations only persist over a scale of order $\xi$. We define the correlation length $l_{c}$ as the distance where $g_{1}(r)$ drops to $1/e$. In Fig. 6(b) we plot the correlation length against $U_{F}$, together with the power law regimes identified.

### v2

@@ I Introduction
A cascade of energy between scales can emerge when energy is injected into a fluid at a particular length scale and dissipated at another scale. This results in the formation of ever smaller structures down to the dissipation scale (direct cascade) [4], or energy accumulating into coherent structures at the system scale (inverse cascade) [17]. Where energy is only transferred between wavenumbers, $k$, of similar magnitude, the cascade is effectively local in momentum-space and power laws can develop in the energy distribution. The slope of power laws can provide clear signatures of particular dynamics, which may otherwise prove opaque. The most well-known example of emergent power-law behavior in fluid turbulence is the Kolmogorov-Obhukov $-5/3$ law [4]. This states that at high Reynolds number in classical incompressible fluids, the distribution of kinetic energy in Fourier-space is given by $E(k)=C\varepsilon^{2/3}k^{-5/3}$, where $C$ is a universal dimensionless constant and $\varepsilon$ is the rate of energy flux per unit mass. Recent experiments [5, 9, 15] and numerical investigations [18, 19] of turbulent BECs have reported power-laws related to predictions from weak-wave turbulence (WWT) theory [20, 21], observed in the wave-action spectrum accessible in experiments.
@@
For weak forcing, we observe $k^{-7/2}$ scaling as reported in experiments and other numerical studies. Further decomposition shows that the quantum pressure is significant throughout the power-law range of $k$, consistent with the role of quantum pressure in short wavelength Bogoliubov quasiparticle excitations [12, 25]. For strong forcing, the system contains many vortices, and approaches a $k^{-7/3}$ power law for the compressible kinetic energy only, consistent with the WWT prediction of inverse particle cascade scaling in the non-condensed regime [19]. The phase coherence length of the GPE wave function drops to the healing length in this regime associated with dense vortex tangles.
@@ II Theoretical Model
Under conditions of weak forcing, the GPE provides a quantitative treatment of the system, directly comparable with experiments. Stronger forcing involves large energy injection and requires careful consideration. Due to the numerical scale of the system and the challenges of running large-scale simulations for very long physical times  [^1] Our physical evolution times of order 10s correspond to around $1300\hbar/\mu$, where $\hbar/\mu$ is the natural unit of time of the homogeneous gas., in this work we take the approach of constructing a minimal GPE model, similar to other recent numerical treatments [5, 23]. Our primary differences are using higher working precision which allows consideration of stronger forcing, and larger spatial sizes in units of the healing length. Ideally, we aim to observe a power-law over a decade of wavenumbers in order to rule out the possibility of falsely attributing power-law behavior to a crossover [36].
@@
where $g=4\pi\hbar^{2}a/m$, $\hbar$ is the reduced Planck constant, $a$ is the $S$-wave scattering length, and $m$ is the atomic mass. For a homogeneous ground state with chemical potential $\mu$ and density $n_{0}=|\psi|^{2}$, we define the healing length $\xi$ via
@@
to coordinates $x-L_{x}$, etc, where $H(x)$ is the Heaviside step. Over a region of width $\ell_{b}=6\xi$ centered on $x=3\xi$ the potential $V_{\textrm{wall}}(x)$ ramps from zero up to $V_{b}=30\mu$, providing sufficiently steep walls for box-confined dynamics while avoiding overly sharp changes in the potential; the shape of the potential causes the flat density to drop away steeply at $x=0$, so that after the shift, the homogeneous region is constrained to ±$\pm L_{x}$. Numerically finding the ground state $\psi_{0}(\mathbf{r})$ of $V_{\textrm{static}}(\mathbf{r})$ using imaginary time evolution for chemical potential $\mu=k_{B}\times 1$nK, results in a total atom number $N\simeq 7.2\times 10^{5}$. The loss of initial anisotropy of the ground state under time evolution provides a useful indicator of the development of steady isotropic turbulence.
@@
This choice ensures that dissipation is only active outside the homogeneous region, $|u|<0.5L_{u}$, and introduces a dissipation scale associated with the escape energy [15]. The momentum required for a particle to escape is a function of the height of the trap walls, and for our trapping potential loss occurs at dimensionless dissipation wavenumber $k_{D}\xi=\sqrt{2V_{b}/\mu}=\sqrt{60}\simeq 7.74$  [^2] This choice also gives sufficient margin between the dissipation $k$-value and the Nyquist $k$-values for a grid of $256^{3}$ points (Appendix A). High-energy particles are thus dissipated well before their spatial grid representation becomes inaccurate. . This energy and spatially selective loss is the only dissipation used in our simulations, and provides the energy sink required for a steady state to form under conditions of steady forcing.
To create turbulence in the forced and damped GPE, we apply $V_{F}$ to the ground state $\psi_{0}$ and evolve until $t=t_{f}\equiv 10$ seconds. This time interval ensures a steady state is reached for all $U_{F}$ values. We use a grid of $256^{3}$ points and time-evolve using a pseudo-spectral Runge-Kutta method, referred to as the RK4IP [39]. A fixed time step $\Delta t=10^{-3}$s is chosen to ensure numerical convergence of all simulations. The code was written in Julia and optimized to achieve a minimal memory footprint and run on GPU hardware using the CUDA.jl library [40, 41]. We achieved a runtime of 38 minutes per second of dynamics on NVIDIA A100 GPU using Float64 representation. For details of the algorithm and error analysis see Appendix A.
Figure 1: Time evolution of density isosurfaces of the cuboid box trap shaken along $z$ with forcing amplitude $U_{F}=\mu$: (a) $t=0.0$s shows the initial ground state of the box trap; (b) $t=0.156$s weak density excitations along $z$; (c) $t=0.391$s nucleation of vortex rings; (d) $t=10.0$s steady-state turbulence. Vortex lines can be seen in red. In the steady state there are many small vortex rings on the surface, while extended long vortex lines are absent from the bulk.
@@ III.1 Vortices and density waves
To outline the general phenomenology, we first present the dynamics that occur at a forcing amplitude of $U_{F}=\mu$. The initial ground state, shown in Fig. 1(a), responds to forcing by weakly oscillating between the sides of the trap, creating weak density waves, as seen in Fig. 1(b). After sufficient energy is injected, the superfluid develops many small vortex rings parallel to the $x-y$ plane [Fig. 1(c)]. After approximately two seconds, the system transitions to a steady state, where the energy per particle oscillates about a steady mean value. The density and vortex distribution becomes isotropic, involving many vortices and significant energy, density, and phase fluctuations [Fig. 1(d)]. For this particular forcing amplitude there are many small vortex rings, but no extended vortex lines.
@@
Fig. 3(a) and (b) show the normalization and energy per particle of the system during excitation. We observe a brief delay between when forcing begins and when particles have sufficient energy to escape the trap, shown in Fig. 3(a). Consequently, the energy per particle of the system rapidly increases [Fig. 3(b)]. Once dissipation takes effect, the energy per particle approaches an approximate steady-state, oscillating about a mean value that increases with $U_{F}$; the system thus approaches an energetic steady state while continuing to lose particles. As in [15], atom loss is approximately linear for weak forcing. In the strongest forcing cases the rate of particle loss decreases over time, resulting in energy per particle gradually increasing.
@@ III.2 Energy decomposition
$E_{\text{kin}}$ $=\frac{m}{2}\int d^{3}\mathbf{r}|\mathbf{u}^{\alpha}(\mathbf{r})|^{2},$ (13)
and $\alpha\in\{i,c,q\}$.
Figure 3: Development of turbulence (a) Atom number over time. The rate of particles escaping the trap increases with $U_{F}$. (b) Total energy per particle over time. (c), (d), (e) Total incompressible, compressible, and quantum pressure kinetic energy during initial period of excitation. (f) Kinetic energy fractions in the steady state ($t=10.0$s) as a function of $U_{F}$. Dashed lines show the four $U_{F}$ values of Fig. 2.
Figures 3(c) and (d) show $E_{\textrm{kin}}^{i}$ and $E_{\textrm{kin}}^{c}$ during the initial period of excitation. Both components oscillate in response to the forcing, acquiring a mean value that increases with $U_{F}$. The kinetic energy components of the steady-state display a nonlinear dependence on $U_{F}$, indicating different regimes of turbulence, shown in Fig. 3(e). The relative importance of compressible and incompressible energy indicates a change from wave to vortex turbulence. Even in the absence of vortices, kinetic energy is stored in density gradients, such as those at the condensate boundary. This accounts for the nonzero $E^{q}_{kin}$ in Fig. 3 for $U_{F}\sim 0$ Figure 3(e) is consistent with the behavior shown in Fig. 2: from $U_{F}=0.2\mu$, $E^{c}_{\textrm{kin}}$ sharply increases as density fluctuations grow [Fig. 3(e)]; for the same $U_{F}$ values, $E^{i}_{\textrm{kin}}$ increases gradually as vortices are present only at the condensate surface. Recent reports of wave-turbulence in BECs generally correspond to these forcing amplitudes ($U_{F}<\mu$) [5, 15, 19, 9], which is consistent with the dominance of compressible kinetic energy in this region.
At $U_{F}\simeq 1.2\mu$, a distinct change is evident in Fig. 3(e). $E^{i}_{\textrm{kin}}$ increases rapidly, associated with the formation of bulk vortices, and $E^{c}_{\textrm{kin}}$ has a much weaker response to increased forcing. Finally, another change in response occurs for $U_{F}>2\mu$, where incompressible kinetic energy and quantum pressure have a weaker response to increased forcing. In Fig. 2 this corresponds to the steady state vortex distribution becoming homogeneous. There may be a limit to the number of vortices that can stably exist within a BEC of a certain size and density, bounding growth of incompressible kinetic energy with forcing. Assuming densely packed vortex rings, the bulk volume divided by the volume of a vortex ring of radius $\sim\pi\xi$, and thickness $\sim 2\xi$, gives an estimated upper bound of $L_{x}L_{y}L_{z}/(2\pi^{3}\xi^{3})\sim 400$ for our system.
This analysis suggests distinct turbulent regimes exist for different forcing strengths. In the following sections, we apply tools of spectral analysis to relate these regimes to previous observations and theoretical predictions, aiming to identify power-law behavior associated with particular regimes in Fig. 3(e).
@@ IV Steady State turbulence
Figure 5: (a)-(c) Decomposed momentum distributions for representative forcing amplitudes of Fig. 2, averaged over one forcing period. (a) As forcing increases a $k^{-3.5}$ scaling regime is evident in both compressible and quantum pressure components. (b) Compressible and quantum pressure components decouple, associated with bulk vortices. Power-law behavior in this region extends only over a very small scale range. (c) For very strong forcing, the compressible wave-action spectrum approaches a $k^{-7/3}$ power-law form. (d) Ratio of total incompressible to compressible kinetic energy, with identified power law regimes. $U_{F}$ values from Fig. 2(b) are shown as dashed lines.
@@
We do not observe any obvious power-law behaviour for $U_{F}\ll\mu$ (not shown). Within a narrow region of forcing, $0.15\mu\lesssim U_{F}\lesssim 0.3\mu$, the compressible density approximates $k^{-3}$. However, the approximate scaling only extends over a factor of 2 in wave number, insufficient to identify power-law behaviour [36]. Starting at $U_{F}=0.4\mu$, we observe a clear $k^{-3.5}$ scaling over nearly a decade of $k\xi$, in both compressible and quantum pressure momentum distributions, shown in Fig. 5(a). This power-law remains intact for a wide interval of forcing energies, $0.4\mu\leq U_{F}\leq 1.2\mu$, and as incompressible energy remains small, the power-law is also evident in the total spectrum $n(k)$. This is consistent with the observed vortex configurations [Fig. 2(a)], where extended bulk vortices remain rare. Despite resembling a power-law, the slope of $n^{i}(k)$ corresponds to a uniform distribution of incompressible energy in $k$-space. In this regime it is notable that $n^{c}(k)\simeq n^{q}(k)$ in the power law region of wavenumbers. We return to this point in Sec. V.1.
@@
For $U_{F}\gtrsim 1.2\mu$, extended vortices develop in the bulk superfluid in the steady-state, and the system enters a transitional regime without clear power-law behavior in any component of kinetic energy [Fig. 5(b)]. The total wave-action spectrum gradually steepens in the ultraviolet due to an increase in the incompressible energy-density, while the compressible and quantum pressure components uncouple, and the amplitude of $n^{c}(k)$ gradually decreases. The appearance of bulk vortices is associated with the breakdown of $k^{-7/2}$ scaling. Eventually, we observe a wide $n^{c}(k)\propto k^{-7/3}$ power-law for strong forcing, $U_{F}\gtrsim 2\mu$, in agreement with the predicted inverse particle cascade for weak forcing in the absence of a condensate [Fig. 5(c)]  [19].
@@
To further understand the different regimes of steady turbulence, in Fig. 5(d) we plot the ratio of incompressible to compressible kinetic energy, with the identified power-law regimes. There is a clear change in the energy dependence on $U_{F}$ at the boundary of each power-law regime. Initially, the ratio is rapidly decreasing, due to compressible excitations paired with a lack of vortex nucleation. With additional energy injected, this is succeeded by a sharp increase in the kinetic energy ratio (as surface vortices begin to proliferate), and the appearance of a $k^{-3.5}$ power-law. The breakdown of $k^{-3.5}$ involves extended vortices entering the bulk in the steady-state, and a steep rise in incompressible energy with $U_{F}$. The system enters a regime of vortex turbulence, but shows no scaling behavior. Finally, for $U_{f}\gtrsim 2\mu$, a $k^{-7/3}$ scaling emerges where vortices are approximately homogeneous and dense, and the incompressible energy responds only weakly to increased forcing. In this regime $E^{i}_{\text{kin}}/E^{c}_{\text{kin}}>1$, a regime we can regard as strong quantum vortex turbulence, in contrast to weak wave turbulence where $E^{i}_{\text{kin}}/E^{c}_{\text{kin}}\ll 1$.
Figure 6: Power law amplitude as a function of forcing strength. $\bar{n}^{c}$ is shown for power laws $n^{c}(k)=\bar{n}^{c}(k\xi)^{-3.5}$ (blue), Fig. 5(a), and $n^{c}(k)=\bar{n}^{c}(k\xi)^{-7/3}$ (red), Fig. 5(c). Error bars express the range of n values for which the power law solution intercepts at least $50\%$ of points in the inertial range. Between the two regimes, the magnitude of $n^{c}(k)$ decays and there is no obvious power law behavior.
@@
For a power law of the form $n(k)=\bar{n}(k\xi)^{-\alpha}$, the amplitude of the power law, $\bar{n}$, can be extracted by plotting a compensated spectra $n(k)(k\xi)^{\alpha}$ (see Fig. 12 for an example of a compensated spectrum). Fig. 6 shows the amplitudes for the two power laws observed in $n^{c}(k)$. The $k^{-7/3}$ regime shows little variation with forcing strength, reflecting the states saturation with densely packed vortices. In contrast, the $k^{-3.5}$ regime demonstrates a proportional increase in the amplitude of the spectrum with forcing.
While Fourier space reveals power-law behavior, we can also examine the corresponding correlation functions in position-space. When power-law behavior weakens, the appearance of characteristic length scales can be more clearly understood in position space. Fourier transformation of the momentum distribution gives the system-averaged spatial two-point correlation of the Gross-Piteavskii wave function, as discussed in Sec. III.3. In Fig. 7(a), we plot the angle-averaged two-point correlation function $g_{1}(r)\equiv G_{\psi\psi}(r)/G_{\psi\psi}(0)$, calculated by transforming $\langle\psi\|\psi\rangle(k)=4\pi k^{2}n(k)$ to position space via (20). The two-point correlator of the ground state ($U_{F}=0$) decays on a scale comparable to the shortest cuboid length $L_{z}$. As $U_{F}$ increases, the length scale is reduced. Eventually, for strong forcing the correlations only persist over a scale of order $\xi$. We define the correlation length $l_{c}$ as the distance where $g_{1}(r)$ drops to $1/e$. In Fig. 7(b) we plot the correlation length against $U_{F}$, together with the power law regimes identified.